OpenStudy (anonymous):

What are the approximate solutions of 4x2 + 3 = −12x to the nearest hundredth? x ≈ −3.23 and x ≈ 0.23 x ≈ −2.72 and x ≈ −0.28 x ≈ 0.28 and x ≈ 2.72 x ≈ −0.23 and x ≈ 3.23

OpenStudy (nurali):

To solve a quadratic equation, put the trinomial on one side and make it equal to 0. Then factor it and make each set of parentheses equal to 0. Then solve both equations to get your solution. 4x² + 3 = -12x +12x = +12x 4x² + 12x + 3 = 0 x = -0.28 , -2.73